OpenStudy (anonymous):

10=6a+7 my answer is a=2 Help

OpenStudy (anonymous):

The answer should be 1/2. If you plug in 2 for a you would get 10=6(2)+7 and 10 =/=19. You would first subtract seven, then divide by 6 and simplify

OpenStudy (anwaarullah):

6a+7=10 6a=10-7 6a=3 a=3/6 a=1/2

OpenStudy (anonymous):

thanks bruh at anwaarullah yours is helpful aswell but he explained it to me first